A freight train, traveling at an average rate of 30 mi/h, leaves san francisco for Los Angeles. 5 hours later, a passenger train, traveling at an average rate of 55 mi/hr, leaves San Francisco for Los Angeles. How far are the trains from San Francisco when the passenger train passes the freight train?
Answer by oberobic(2304)

Freight train:
rate = 30 mph
time = t
.
Passenger train:
rate = 55 mph
time = t-5
.
When the passenger train overtakes the freight train, they will have traveled the same distance.
Therefore,
.
30*t = 55*(t-5)
.
30t = 55t - 275
.
-25t = -275
.
t = 11
.
In 11 hours the passenger train will overtake the freight train.
.
Always check the answer to be sure it is correct.
.
30*11 = 330 miles traveled by the freight train
55*(11-5)= 55*6 = 330 miles traveled by the passenger train
.
Done.